OpenStudy (anonymous):

subtract and simplify: (2)/(3x+5)-(1)/(6x+10)

OpenStudy (kropot72):

Using (3x - 5)(6x + 10) as a common denoninator:\[((2\times(6x+10))-(3x-5))\div((3x-5)\times(6x+10))\] This reduces to: \[(9x+25)\div((3x-5)\times(6x+10))\]